OpenStudy (anonymous):

Sum of a Geometric Series Hey, can I please have some help with this question. Evaluate: 18 + 9 + 4 1/2 + ... + 9/64 So A = 18 R = 2 How do I find "N"?

OpenStudy (shamil98):

N = the number of terms ..

OpenStudy (shamil98):

a1 = 18 r = 2 your ration is 1/2

OpenStudy (shamil98):

ratio of terms

OpenStudy (shamil98):

18/2 = 9 9/2 = 4 1/2

OpenStudy (anonymous):

Yeah, I got it to 18x2^n-1 = 9/64

OpenStudy (anonymous):

Then I went 2^n-1 = 81/32 but I do not know where to go from there

OpenStudy (anonymous):

I do not know how to get it to same base

OpenStudy (shamil98):

take the log

OpenStudy (anonymous):

I do not know how to do that, we have not done logs yet

OpenStudy (anonymous):

Just realised I had ratio of 2 instead of 1/2 Hopefully I can work it out now

OpenStudy (anonymous):

Nope still have not got it

OpenStudy (shamil98):

just keep dividing the ratios till you 9/64.. a1 = 18 a2 = 9 a3 = 4.5 a4 = 2.25 a 5 = 1.125 a6 = .5625 a7 = .28125 a8 = 0.140625 and there you have your answer n = 8 9/64 = 0 .140625

OpenStudy (anonymous):

So no easier way? Just have to keep dividing?

OpenStudy (nincompoop):

there's an easier way... if you read what I gave you

OpenStudy (shamil98):

well there probably is. i'm just too tired to find what it is.

OpenStudy (anonymous):

Fair enough Shamil And nin this type of question was not there

OpenStudy (anonymous):

Dinner :P back in a bit

OpenStudy (nincompoop):

omg... you want to be spoon-fed

OpenStudy (anonymous):

Yeah because I can not understand the question so i need explaining

OpenStudy (anonymous):

Any help?

OpenStudy (anonymous):

Don't worry, I had times 18 instead of divided 18 Thanks for the help